OpenStudy (anonymous):

I need help understanding how to find the domain and range algebraically and graphically..

OpenStudy (anonymous):

an example would be nice

OpenStudy (anonymous):

if you have the graph it is more or less easy you can look to see what possible \(x\) values you can use and also the possible \(y\) values

OpenStudy (anonymous):

I have two: y=\[\sqrt{x-4}\] y=\[\sqrt{x ^{2}-4}\]

OpenStudy (anonymous):

It takes light to roughly 3.0 x 10–9 seconds to travel one meter. Write this time in standard form.need some help

OpenStudy (anonymous):

in both of these you have to make sure the expresson inside the radical is not negative, so set \(x-4\geq0\) and solve for \(x\) (in one easy step) for the first one that will give you the domain

OpenStudy (anonymous):

the range is \([0,\infty)\) because the square root means the positive square root

OpenStudy (anonymous):

second one is harder but not really you have \(y=\sqrt{x^2-4}\) so again \(x^2-4\geq 0\) is what you need \(x^2-4\) is a parabola that opens up with zeros at \(-2\) and \(2\) so it will be negative between the zeros and positive outside them therefore the domain is \((-\infty.-2]\cup [2,\infty)\)

OpenStudy (anonymous):

thanks!